OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -9 - 3t. Find the instantaneous velocity at t = 8 by finding the derivative. @e.mccormick @jim_thompson5910 would this sort of be like the other ones? the same process?

OpenStudy (e.mccormick):

Yes. Instant velocity is found using the derivative.

OpenStudy (anonymous):

i got -3

OpenStudy (e.mccormick):

Yes. When there is no power, it is pretty easy that way.

OpenStudy (anonymous):

so thats it? or

OpenStudy (anonymous):

@e.mccormick

OpenStudy (e.mccormick):

That is it.

OpenStudy (anonymous):

what about the whole t=8 part so i ignore it?

OpenStudy (e.mccormick):

It has a constant velocity, so it is the same at any point.

OpenStudy (anonymous):

so if i turn this is i will be correct: d/dx(-9-3x)=-3 -d/dx(9)-d/dx(3x) d/dx=(9)=0 d/dx(3x0=3 =0-3 =-3

OpenStudy (e.mccormick):

If it was \(s(t) = -9 - 3t^2\) then the velocity would be changing and there would be more to do. But for \(s(t) = -9 - 3t\) it is just -3.

OpenStudy (e.mccormick):

Yah.

OpenStudy (anonymous):

thank you c: